# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

### v1

@@ 3.4.2 Spatio-Temporal Representation via Customized LSTM
To effectively model both spatial and temporal dependencies within a video, frame-level spatial features $\mathbf{F}_{t}\in\mathbb{R}^{1024}$ are first extracted for each of the $T$ frames, forming a sequential input $\mathbf{F}=[\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{T}]$ as defined in Equation 10. These features are then fed into a customized LSTM network, which learns temporal transitions across time steps. At each time step $t$, the LSTM computes the input gate $\mathbf{i}_{t}$, forget gate $\mathbf{f}_{t}$, output gate $\mathbf{o}_{t}$, and the candidate cell state $\tilde{\mathbf{c}}t$ based on the previous hidden state $\mathbf{h}{t-1}$ and the current input $\mathbf{F}_{t}$, as shown in Equation 11. The cell state $\mathbf{c}_{t}$ is updated using Equation 12, and the corresponding hidden state $\mathbf{h}_{t}$ is computed according to Equation 13.
@@ 3.5 Features Selection
3.5.1 Dynamic Fitness Sharing
@@ 3.5.2 Proposed Novel Ensemble Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
An ensemble fitness function is proposed for feature set optimization in HAR. The proposed method integrates multiple criteria, classification accuracy, feature redundancy, uniqueness, and complexity into a unified optimization framework using a dynamically adaptive weighting strategy. This ensemble fitness function is designed to guide the give GA-based search process to identify an optimal subset of features that are both discriminative and compact.
@@
3.5.3 Mathematical Interpretation of our Proposed ADFSA Approach
The proposed approach combines multiple factors, accuracy, redundancy, uniqueness-boost, and complexity into a dynamic FS framework. The following is the mathematical representation of the ADFSA scheme incorporated in the proposed methodology.
@@ 3.5.4 Key Advantages of the Proposed Framework
The proposed approach introduces a comprehensive and adaptive FS framework that integrates four fundamental aspects, accuracy, redundancy, uniqueness, and complexity into a unified fitness function. This multi-objective formulation ensures a balanced trade-off between classification performance and model simplicity. One of the key innovations lies in the dynamic adjustment of the weights associated with each fitness component during the evolutionary process. This adaptive mechanism fosters exploration in the initial generations and shifts towards exploitation in later stages, which enhances convergence and helps avoid premature stagnation in local optima. Additionally, the method incorporates a redundancy penalty based on mutual information, effectively minimizing the selection of highly correlated features and ensuring that only the most informative attributes are retained. To further improve population diversity, a uniqueness term is introduced that rewards diverse feature subsets across generations, increasing the likelihood of identifying globally optimal solutions. The inclusion of a complexity component drives the selection of smaller, more efficient feature sets, thereby reducing the computational cost and facilitating faster model inference, an essential requirement for real-time and embedded HAR systems. Empirical evaluations demonstrate that the proposed method outperforms traditional FS techniques such as PCA and RFE in terms of classification accuracy, even with fewer selected features. Moreover, the approach is classifier-agnostic and compatible with a variety of learning algorithms, including SVM, KNN, Random Forest (RF), and Logistic Regression (LogReg). Its scalability makes it particularly suitable for handling high-dimensional feature vectors derived from deep learning models, addressing the limitations of conventional methods related to scalability and interpretability. Overall, the proposed ensemble-based, dynamically adaptive, and diversity-aware framework presents a novel and effective solution for FS in HAR and other pattern recognition domains.
@@
In our proposed approach, redundancy and uniqueness-boost mechanisms ensure that the algorithm does not converge too early on a suboptimal feature subset. Furthermore, the inclusion of model complexity into the fitness function will produce simpler models, which are more computationally efficient and easier to interpret. Incorporating flexibility across different types of FS tasks, and penalty terms having adaptive natures, allows fine-tuning based on the specific problem or dataset.
3.6 Lightweight Machine Learning Classifiers
@@
To assess the effectiveness of the proposed FS method, we perform a comparative evaluation against two widely used techniques: PCA and RFE. PCA is a linear dimensionality reduction approach that transforms the feature space into orthogonal components capturing the maximum variance. RFE, on the other hand, iteratively removes the least important features based on model weights, gradually selecting the most significant subset.
@@
Our proposed method employs a custom-designed fitness function that adaptively balances classification accuracy $f_{acc}$, feature redundancy $f_{red}$, uniqueness $f_{uni}$, and complexity $f_{comp}$. The dynamic weighting mechanism evolves over generations, initially favoring diversity and redundancy control, and later shifting focus toward accuracy and compactness.
@@
The optimized feature vector $\mathbf{f}^{*}$, generated through the proposed GA-based FS framework, is evaluated using multiple lightweight machine learning classifiers to assess its classification capability and generalizability. The SVM, as formulated in Equation 28, determines the class label by computing the sign of a weighted linear combination of features and bias, effectively creating a maximum-margin decision boundary. The RF classifier, shown in Equation 29, aggregates predictions from an ensemble of $T$ decision trees and outputs the majority-voted label, thereby reducing variance and enhancing robustness. The k-Nearest Neighbors (k-NN) approach, expressed in Equation 30, relies on the most frequent class among the $k$ closest samples in the feature space, making it highly adaptive to local patterns in the data. LogReg, represented in Equation 31, calculates the sigmoid-activated probability and assigns the class label based on a threshold decision rule. Together, these classifiers provide a diverse set of perspectives to validate the quality of the selected features. To further benchmark the effectiveness of the proposed FS method, classical techniques like PCA and RFE are also applied under identical evaluation settings. All models are evaluated in terms of both classification accuracy and inference time to validate the practical efficiency of the selected features.
@@ 4.1 Dataset
The YouTube Action Dataset [11] is a widely utilized benchmark in the field of human activity recognition. It comprises a diverse collection of videos sourced from YouTube, capturing real-world human actions across a variety of contexts and environments. The dataset includes multiple action categories such as walking, running, biking, and hand-waving, among others, providing a challenging testbed for evaluating action recognition algorithms.
@@
4.2 Implementation
Overall the research was implemented into two different phases, in the first phase, spatio-temporal features were extracted using a customized Inception-V3 and LSTM network, and in the second phase, the extracted features were passed through the proposed ADFSA FS method to obtain compact and robust features without affecting the performance of the underlying classifier.
@@ 4.2.1 Spatial Feature Extraction Phase
This work employs a deep CNN architecture to extract discriminative features from video frames. Extensive experiments were carried out by training several deep CNN models, as summarized in Table 3. Among the evaluated architectures, InceptionV3 consistently outperformed the others, demonstrating good recognition accuracy. Trainng of selected networks and overall proposed framework is presented in Figure 6. Consequently, it was selected and customized for the HAR task, capturing diverse and robust features region and edge, leading to enhanced performance. The training and loss curves of the optimized model are illustrated in Figures 7(a) and 7(b), respectively.
@@
4.2.2 Temporal Feature Learning using LSTM
To capture the temporal dynamics present across consecutive video frames, an LSTM network is employed. Features of 20 frames extracted from customized Inception-V3 in a sequential manner are fed into the LSTM, learning long-range dependencies between consecutive individual frames, retaining historical context over time. In this setup, the LSTM network consists of 128 memory units and is trained using the frame-wise feature vectors in sequence. This enables the model to understand motion patterns, temporal transitions, and dependencies between frames that are critical for activity recognition. After training, the LSTM outputs a feature representation that integrates both spatial information (from InceptionV3) and temporal dependencies (learned via LSTM). These fused spatio-temporal features are then used for subsequent feature optimization and classification tasks.
@@ 4.2.2 Temporal Feature Learning using LSTM
Customized Inception-V3 LSTM 99.32%
@@ 4.2.3 Features Selection via ADFSA
To enhance the quality of the extracted features, an optimization step was carried out using the proposed ADFSA algorithm. This method leverages a GA guided by an ensemble-based fitness function that integrates four complementary evaluation criteria. First, classification accuracy is assessed using stratified 5-fold cross-validation with a linear SVM, ensuring robust performance estimation. Second, a redundancy penalty based on mutual information is incorporated to discourage the selection of highly correlated features, promoting a more informative subset. Third, a uniqueness component is introduced to enhance population diversity by penalizing the recurrence of identical feature subsets across generations. Finally, a complexity term penalizes larger subsets, thereby encouraging compact and computationally efficient FSs. Notably, the relative weights of these components are adaptively updated throughout the evolutionary process, guided by the normalized progression of the population’s fitness. This dynamic adjustment ensures a balance between exploration and exploitation, ultimately leading to the selection of high-quality, low-redundancy feature sets well-suited for classification tasks. To assess the effectiveness of the proposed method, two widely used FS techniques were implemented, PCA [25] and RFE [26]. The effectiveness of various selection methods was assessed using a range of classification algorithms, comprising SVM, RF, KNN, and LogReg.
5 Results and Discussions
Extensive experiments were carried out using HAR datasets to assess the effectiveness of the proposed ensemble dynamic fitness function for FS. A comparison was made between the proposed method and two established techniques, PCA and RFE. The findings show that the proposed method surpasses both PCA and RFE in terms of classification accuracy, feature compactness, and inference efficiency. By dynamically adjusting the contributions of various components, such as accuracy, redundancy, uniqueness, and complexity, the fitness function guides the evolutionary process toward more optimal feature subsets. This adaptive approach led to the selection of significantly fewer features (as few as 7 out of 128) without compromising, and often enhancing, the model’s predictive performance. The top-performing setup achieved a test accuracy of 99.58% using only 7 features, while also reducing average inference time. These results confirm the efficacy of the dynamic ensemble strategy in identifying informative and concise feature representations for HAR, positioning it as a promising option for real-time and resource-limited applications.
As shown in Figure 1, the extracted spatial features are subsequently fed into LSTM network comprising 128 LSTM units. Given that video data consists of sequential frames exhibiting temporal continuity, the LSTM network is well-suited to capture and learn these temporal dependencies. Once the LSTM model is trained, temporal features are extracted by passing the input sequences through the network. This process results in a feature matrix that effectively integrates both spatial and temporal characteristics of the video data.
5.1 2D Features Space Visualization
Visualization shown in Figure 8(a) appears to show a relatively linear, diagonal clustering trend across the latent feature space. The color gradient representing ”Feature Labels” is moderately well-separated, with neighboring clusters transitioning smoothly. However, there is still a significant degree of overlap between different class regions, especially in the bottom-left and center areas. This overlap may suggest that while LSTM has captured some degree of class-specific information, it lacks sharp class boundary definitions, potentially due to noise or insufficient feature discrimination.
The RFE approach exhibits a more refined clustering compared to LSTM as shown in Figure 8(b). The class transitions (as per the color bar) are smoother, and the data points are more tightly grouped. However, there is still observable overlap at the cluster borders. RFE seems to preserve the intrinsic feature structure more effectively than LSTM, indicating better dimensional reduction and selection fidelity. Nevertheless, this method may still struggle with high inter-class similarity, particularly in the mid-spectrum of the feature labels.
(a) Using no FS
(b) Using RFE
Figure 8: Feature space visualization with no FS vs RFE.
(a) Using PCA
(b) Using (proposed) ADFSA
Figure 9: Feature space visualization with PCA vs Proposed ADFSA.
The most notable contrast is seen in the visualization using the proposed ADFSA method shown in Figure 9(b). This plot shows distinct and well-separated clusters, with a clearer demarcation among different feature label zones. The visualized feature space demonstrates improved intra-class compactness and inter-class separation, indicating superior FS and reduction capabilities. The spread of clusters in multiple directions suggests that ADFSA captures more complex relationships within the data, which may translate to better generalization and classification performance in downstream tasks.
5.2 Evaluation Metrics
To comprehensively evaluate the performance of the proposed framework, several quantitative metrics were employed. Classification accuracy (%) was computed on the test set to assess the overall predictive capability of the model by Equation 32. In addition, the precision, recall, F1-score, were used to provide a more balanced evaluation, particularly in the presence of class imbalance by Equations 3335. The number of selected features was recorded to reflect the dimensionality reduction achieved through the FS process. To assess computational efficiency, the average inference time (in milliseconds) was measured over five independent runs. Furthermore, memory usage (in KB) of the selected features was analyzed to evaluate the memory footprint of the model. These metrics together offer a holistic view of the classification performance, computational efficiency, and practicality of the proposed approach.
$\mathrm{Accuracy}=\frac{TP}{TP+FP+FN}$ (32)
$\mathrm{Precision}=\frac{TP}{TP+FP}$ (33)
$\mathrm{Recall}=\frac{TP}{TP+FN}$ (34)
@@ 5.3 Hardware and Implementation
All machine learning classifiers and FS algorithms were implemented and executed using the Spyder integrated development environment (IDE).The codebase was implemented in python 3.10, utilizing TensorFlow/Keras, Scikit-learn, openCV, and NumPy libraries. The experiments related to traditional ML models, such as SVM, RF, KNN, and LogReg, were conducted on a system equipped with an Intel Core i7 10th-generation processor and 16 GB of RAM. For the training of deep CNN, including InceptionV3 and LSTM-based models, a high-performance computing setup was used, consisting of an HP desktop powered by an Intel Core-i9 9th-generation processor, 64 GB of RAM, and an NVIDIA GeForce RTX 2080 Ti GPU, all the cofiguration is given in Table 4. This dual-platform configuration was essential to accommodate both lightweight ML tasks and compute-intensive deep learning workloads effectively.
@@
Temporal Model Customized LSTM with 128 units

### v2

@@ 3.2.2 Spatio-Temporal Representation via Attention-Augmented LSTM
To effectively model both spatial and temporal dependencies within a video, frame-level spatial features $\mathbf{F}_{t}\in\mathbb{R}^{1024}$ are first extracted for each of the $T$ frames, forming a sequential input $\mathbf{F}=[\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{T}]$ as defined in Equation 10. These features are then fed into a customized AA-LSTM network, which learns temporal transitions across time steps. At each time step $t$, the AA-LSTM computes the input gate $\mathbf{i}_{t}$, forget gate $\mathbf{f}_{t}$, output gate $\mathbf{o}_{t}$, and the candidate cell state $\tilde{\mathbf{c}}t$ based on the previous hidden state $\mathbf{h}{t-1}$ and the current input $\mathbf{F}_{t}$, as shown in Equation 11. The cell state $\mathbf{c}_{t}$ is updated using Equation 12, and the corresponding hidden state $\mathbf{h}_{t}$ is computed according to Equation 13.
@@ 3.3 Features Selection
3.3.1 Dynamic Fitness Sharing for Feature Selection
@@ 3.3.2 Proposed Novel Ensemble Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
An ensemble fitness function is proposed for feature set optimization in HAR. The proposed technique integrates multiple criteria, classification accuracy, feature redundancy, uniqueness, and complexity into a unified optimization framework using a dynamically adaptive weighting strategy. This ensemble fitness function is designed to guide the give GA-based search process to identify an optimal subset of features that are both discriminative and compact.
@@
3.3.3 Mathematical Interpretation of our Proposed ADFSA Technique
The proposed technique combines multiple factors, accuracy, redundancy, uniqueness-boost, and complexity into a dynamic FS framework. The following is the mathematical representation of the ADFSA scheme incorporated in the proposed framework.
@@ 3.3.4 Key Advantages of the Proposed Framework
The proposed framework introduces a comprehensive and adaptive FS technique that integrates four fundamental aspects, accuracy, redundancy, uniqueness, and complexity into a unified fitness function. This multi-objective formulation ensures a balanced trade-off between classification performance and model simplicity. One of the key innovations lies in the dynamic adjustment of the weights associated with each fitness component during the evolutionary process. This adaptive mechanism fosters exploration in the initial generations and shifts towards exploitation in later stages, which enhances convergence and helps avoid premature stagnation in local optima. Additionally, the method incorporates a redundancy penalty based on mutual information, effectively minimizing the selection of highly correlated features and ensuring that only the most informative attributes are retained. To further improve population diversity, a uniqueness term is introduced that rewards diverse feature subsets across generations, increasing the likelihood of identifying globally optimal solutions. The inclusion of a complexity component drives the selection of smaller, more efficient feature sets, thereby reducing the computational cost and facilitating faster model inference, an essential requirement for real-time and embedded HAR systems. Empirical evaluations demonstrate that the proposed technique outperforms traditional FS techniques such as PCA and RFE in terms of classification accuracy, even with fewer selected features. Moreover, the approach is classifier-agnostic and compatible with a variety of learning algorithms, including SVM, KNN, Random Forest (RF), and Logistic Regression (LogReg). Its scalability makes it particularly suitable for handling high-dimensional feature vectors derived from deep learning models, addressing the limitations of conventional methods related to scalability and interpretability. Overall, the proposed ensemble-based, dynamically adaptive, and diversity-aware technique presents a novel and effective solution for FS in HAR and other pattern recognition domains.
@@
In our proposed technique, redundancy and uniqueness-boost mechanisms ensure that the algorithm does not converge too early on a suboptimal feature subset. Furthermore, the inclusion of model complexity into the fitness function will produce simpler models, which are more computationally efficient and easier to interpret. Incorporating flexibility across different types of FS tasks, and penalty terms having adaptive natures, allows fine-tuning based on the specific problem or dataset.
3.4 Classification
@@
To assess the effectiveness of the proposed FS technique, we perform a comparative evaluation against two widely used techniques: PCA and RFE. PCA is a linear dimensionality reduction approach that transforms the feature space into orthogonal components capturing the maximum variance. RFE, on the other hand, iteratively removes the least important features based on model weights, gradually selecting the most significant subset.
@@
Our proposed technique employs a custom-designed fitness function that adaptively balances classification accuracy $f_{acc}$, feature redundancy $f_{red}$, uniqueness $f_{uni}$, and complexity $f_{comp}$. The dynamic weighting mechanism evolves over generations, initially favoring diversity and redundancy control, and later shifting focus toward accuracy and compactness.
@@
The optimized feature vector $\mathbf{f}^{*}$, generated through the proposed GA-based FS technique, is evaluated using multiple lightweight machine learning classifiers to assess its classification capability and generalizability. The SVM, as formulated in Equation 28, determines the class label by computing the sign of a weighted linear combination of features and bias, effectively creating a maximum-margin decision boundary. The RF classifier, shown in Equation 29, aggregates predictions from an ensemble of $T$ decision trees and outputs the majority-voted label, thereby reducing variance and enhancing robustness. The k-Nearest Neighbors (k-NN) approach, expressed in Equation 30, relies on the most frequent class among the $k$ closest samples in the feature space, making it highly adaptive to local patterns in the data. LogReg, represented in Equation 31, calculates the sigmoid-activated probability and assigns the class label based on a threshold decision rule. Together, these classifiers provide a diverse set of perspectives to validate the quality of the selected features. To further benchmark the effectiveness of the proposed FS technique, classical techniques like PCA and RFE are also applied under identical evaluation settings. All models are evaluated in terms of both classification accuracy and inference time to validate the practical efficiency of the selected features.
@@ 4 Experimental Setup
The experimental setup of this research is structured into four main components. First, the dataset is introduced, highlighting its composition and class distribution. Second, a two-phase framework is applied, where spatio-temporal features are extracted using a customized deep learning architecture and then refined through the proposed feature selection technique. Third, the hardware and software environment is outlined to specify the implementation platform and computational resources used. Together, these components provide a clear foundation for evaluating the effectiveness and efficiency of the proposed framework.
@@ 4.1 Dataset
The YouTube Action Dataset [11] is a widely used benchmark in human activity recognition, consisting of diverse videos collected from YouTube that capture real-world human actions in varied contexts and environments. It covers multiple action categories such as walking, running, biking, and hand-waving, among others, making it a challenging testbed for evaluating recognition algorithms. In this research, the dataset was divided into training and testing sets using an $80/20$ split to ensure balanced evaluation of the proposed framework.
@@
4.2 Two-Phase Feature Extraction and Selection Framework
Overall the research was implemented into two different phases. In the first phase, spatio-temporal features were extracted using a customized Inception-V3 and AA-LSTM network, and in the second phase, the extracted features were passed through the proposed ADFSA FS technique to obtain compact and robust features without affecting the performance of the underlying classifier.
@@ 4.2.1 Spatial Feature Extraction Phase
This work employs a deep CNN architecture to extract discriminative features from video frames. Extensive experiments were carried out by training several deep CNN models, as summarized in Table 3. Among the evaluated architectures, InceptionV3 consistently outperformed the others, demonstrating good recognition accuracy. Trainng of selected networks and the overall proposed framework is presented in Figure 6. Consequently, it was selected and customized for the HAR task, capturing diverse and robust features region and edge, leading to enhanced performance. The training and loss curves of the optimized model are illustrated in Figures 7(a) and 7(b), respectively.
@@
4.2.2 Temporal Feature Learning using AA-LSTM
To capture the temporal dynamics present across consecutive video frames, an AA-LSTM network is employed. Features of 20 frames extracted from customized Inception-V3 in a sequential manner are fed into the AA-LSTM, learning long-range dependencies between consecutive individual frames, retaining historical context over time. In this setup, the AA-LSTM network consists of 128 memory units and is trained using the frame-wise feature vectors in sequence. This enables the model to understand motion patterns, temporal transitions, and dependencies between frames that are critical for activity recognition. After training, the AA-LSTM outputs a feature representation that integrates both spatial information (from InceptionV3) and temporal dependencies (learned via AA-LSTM). These fused spatio-temporal features are then used for subsequent feature optimization and classification tasks.
@@ 4.2.2 Temporal Feature Learning using AA-LSTM
Customized Inception-V3 + AA-LSTM 99.32%
@@ 4.2.3 Features Selection via ADFSA
To enhance the quality of the extracted features, an optimization step was carried out using the proposed ADFSA algorithm. This method leverages a GA guided by an ensemble-based fitness function that integrates four complementary evaluation criteria. First, classification accuracy is assessed using stratified 5-fold cross-validation with a linear SVM, ensuring robust performance estimation. Second, a redundancy penalty based on mutual information is incorporated to discourage the selection of highly correlated features, promoting a more informative subset. Third, a uniqueness component is introduced to enhance population diversity by penalizing the recurrence of identical feature subsets across generations. Finally, a complexity term penalizes larger subsets, thereby encouraging compact and computationally efficient FSs. Notably, the relative weights of these components are adaptively updated throughout the evolutionary process, guided by the normalized progression of the population’s fitness. This dynamic adjustment ensures a balance between exploration and exploitation, ultimately leading to the selection of high-quality, low-redundancy feature sets well-suited for classification tasks. To assess the effectiveness of the proposed technique, two widely used FS techniques were implemented, PCA [25] and RFE [26]. The effectiveness of various selection methods was assessed using a range of classification algorithms, comprising SVM, RF, KNN, and LogReg.
@@ 4.3 Hardware and Implementation
All machine learning classifiers and FS algorithms were implemented and executed using the Spyder integrated development environment (IDE).The codebase was implemented in python 3.10, utilizing TensorFlow/Keras, Scikit-learn, openCV, and NumPy libraries. The experiments related to traditional ML models, such as SVM, RF, KNN, and LogReg, were conducted on a system equipped with an Intel Core i7 10th-generation processor and 16 GB of RAM. For the training of deep CNN, including customized InceptionV3 and AA-LSTM-based models, a high-performance computing setup was used, consisting of an HP desktop powered by an Intel Core-i9 9th-generation processor, 64 GB of RAM, and an NVIDIA GeForce RTX 2080 Ti GPU, all the cofiguration is given in Table 4. This dual-platform configuration was essential to accommodate both lightweight ML tasks and compute-intensive deep learning workloads effectively.
@@
Temporal Model Customized AA-LSTM with 128 units
@@
5 Results and Discussions
Extensive experiments were conducted on benchmark HAR datasets to evaluate the proposed dynamic composite fitness function for feature selection. Compared with PCA and RFE, the method achieved higher accuracy, selected fewer features, and improved inference speed. By adjusting the balance between accuracy, redundancy, uniqueness, and complexity, the fitness function guided the search toward compact yet effective feature subsets. In one case, only 7 out of 128 features were selected while still reaching a test accuracy of 99.58%, along with reduced memory use and faster execution. These results show that the method can provide accurate and efficient recognition, even in real-time or resource-limited settings.
The detailed results are presented in the following subsections. First, 2D feature space visualizations are shown to highlight the separability of the optimized features. Next, evaluation metrics and implementation settings are described. This is followed by a performance summary along with memory and training time analysis. Additional measures such as F1-score, precision, and recall are then reported to provide a broader view of classification performance. An ablation study is included to examine the contribution of each component of the fitness function. Finally, the proposed technique is compared with state-of-the-art approaches to demonstrate its overall effectiveness.
5.1 2D Features Space Visualization
Visualization shown in Figure 8(a) appears to show a relatively linear, diagonal clustering trend across the latent feature space. The color gradient representing ”Feature Labels” is moderately well-separated, with neighboring clusters transitioning smoothly. However, there is still a significant degree of overlap between different class regions, especially in the bottom-left and center areas. This overlap may suggest that while AA-LSTM has captured some degree of class-specific information, it lacks sharp class boundary definitions, potentially due to noise or insufficient feature discrimination.
The RFE approach exhibits a more refined clustering compared to AA-LSTM as shown in Figure 8(b). The class transitions (as per the color bar) are smoother, and the data points are more tightly grouped. However, there is still observable overlap at the cluster borders. RFE seems to preserve the intrinsic feature structure more effectively than AA-LSTM, indicating better dimensional reduction and selection fidelity. Nevertheless, this method may still struggle with high inter-class similarity, particularly in the mid-spectrum of the feature labels.
(a) Using no FS
(b) Using RFE
Figure 8: Feature space visualization with no FS vs RFE.
(a) Using PCA
(b) Using (proposed) ADFSA
Figure 9: Feature space visualization with PCA vs Proposed ADFSA.
The most notable contrast is seen in the visualization using the proposed ADFSA technique shown in Figure 9(b). This plot shows distinct and well-separated clusters, with a clearer demarcation among different feature label zones. The visualized feature space demonstrates improved intra-class compactness and inter-class separation, indicating superior FS and reduction capabilities. The spread of clusters in multiple directions suggests that ADFSA captures more complex relationships within the data, which may translate to better generalization and classification performance in downstream tasks.
5.2 Evaluation Metrics
To comprehensively evaluate the performance of the proposed framework, several quantitative metrics were employed. Classification accuracy (%) was computed on the test set to assess the overall predictive capability of the model by Equation 32. In addition, the precision, recall, F1-score, were used to provide a more balanced evaluation, particularly in the presence of class imbalance by Equations 3335. The number of selected features was recorded to reflect the dimensionality reduction achieved through the FS process. To assess computational efficiency, the average inference time (in milliseconds) was measured over five independent runs. Furthermore, memory usage (in KB) of the selected features was analyzed to evaluate the memory footprint of the model. These metrics together offer a holistic view of the classification performance, computational efficiency, and practicality of the proposed framework.
$\mathrm{Accuracy}=\frac{TP}{TP+FP+FN}$ (32)
$\mathrm{Precision}=\frac{TP}{TP+FP}$ (33)
$\mathrm{Recall}=\frac{TP}{TP+FN}$ (34)